OpenStudy (anonymous):

what is the answer ? solve for x:|x+3|=8 ?

OpenStudy (cherrilyn):

x = 5 or x = -11

OpenStudy (anonymous):

x=5,-11

OpenStudy (anonymous):

|x+3|=8 =>x+3=8 or x+3=-8, so you get it.

OpenStudy (anonymous):

x+3=8 x+3=-8 x=8-3 x= -8-3 x = 5 x = -11